The largest number of 100 W electric light bulbs which can be operated from a 240 V supply fitted with a 13 A fuse is


1
Expert's answer
2021-08-26T08:11:05-0400

Given:

"V=240\:\rm V"

"I_{\rm max}=13\:\rm A"

"P=100\:\rm W"


The maximum power in circuit

"P_{\rm max}=I_{\rm max}V=13*240=3120\:\rm W"

Hence, the largest number of 100 W electric light bulbs which can be operated in circuit

"N=\frac{P_{\rm max}}{P}=\frac{3120}{100}=31"